twelve divided by the sum of x and 2 equals the quotient of 4 and the difference of x and 2 find x
Answer by stanbon(48516)

twelve divided by the sum of x and 2 equals the quotient of 4 and the difference of x and 2 find x
------
12/(x+2) = 4/(x-2)
Cross-multiply to get:
12x-24 = 4x+8
8x = 32
x = 4
===========
Cheers,
Stan H.